OpenStudy (kellyspeakslouder):

f(x) = Square root of quantity x plus nine.; g(x) = 8x - 13 Find f(g(x)). A. f(g(x)) = 2Square root of quantity two x minus one. B. f(g(x)) = 8Square root of quantity x minus four. C. f(g(x)) = 2Square root of quantity two x plus one. D. f(g(x)) = 8Square root of quantity x plus nine. - 13

OpenStudy (kellyspeakslouder):

@zepdrix

zepdrix (zepdrix):

AHHHH quiet down kellbell >.< too loud! Anyway, so we've got some stuff listed,\[\large\rm f=\sqrt{x+9}\qquad\qquad\qquad g=8x-13\]Our problem is asking us to find the composition, g inside of f.\[\large\rm f(g)=\sqrt{g+9}\]So we take the entire function g, and replace it with every x in f.

OpenStudy (kellyspeakslouder):

lol! sorry, this is just so confusing to me. I will try my best though..

zepdrix (zepdrix):

\[\large\rm f=\sqrt{\color{orangered}{x}+9}\]When we take a composition, we replace all of the x with an `entire function`.\[\large\rm f(\color{orangered}{g})=\sqrt{\color{orangered}{g}+9}\]On the right side of the equals let's plug in our g, ya?\[\large\rm f(\color{orangered}{g})=\sqrt{\color{orangered}{8x-13}+9}\]Confusing? What do you think KellySpaghetti?

OpenStudy (kellyspeakslouder):

it would be f(g)= sqrt 8g minus 13 + 9, right?..

zepdrix (zepdrix):

No no :O I already put the g inside. Don't put g inside of the g. That's crazy talk.

OpenStudy (kellyspeakslouder):

oh goodness lol okay

zepdrix (zepdrix):

We started with f. We plugged the letter g in place of x. We then replaced that g with its equation. Now simplify from that point,\[\large\rm f(\color{orangered}{g})=\sqrt{\color{orangered}{8x-13}+9}\]

zepdrix (zepdrix):

No more fancy plugging. Just simplify.

OpenStudy (kellyspeakslouder):

this might take me a while lol

zepdrix (zepdrix):

Your hair is ridiculous kelly -_- I'm telling your mom

zepdrix (zepdrix):

Just combine -13 and 9. And then we'll have one tiny step after that :)

OpenStudy (kellyspeakslouder):

negative 4 and nooooo don't tell her </3

zepdrix (zepdrix):

\[\large\rm f(g)=\sqrt{8x-4}\]Ok great. These terms have something in common that we can factor out.

OpenStudy (kellyspeakslouder):

4 goes into both of them? or is it 2 that goes into both of them?

zepdrix (zepdrix):

4 seems good

OpenStudy (kellyspeakslouder):

awesomesauce

OpenStudy (kellyspeakslouder):

would the answer be D?

OpenStudy (kellyspeakslouder):

hellow?..

zepdrix (zepdrix):

Oh sorry you were taking a little while so I went and made a sandwich :D and then read a book :D and watched a full length movie.

OpenStudy (kellyspeakslouder):

nice! what movie?

zepdrix (zepdrix):

LOL I was just kidding :D

zepdrix (zepdrix):

So you pull the 4 out of each value,\[\large\rm f(g)=\sqrt{4(2x-1)}\]

zepdrix (zepdrix):

Then you pull the 4 out of the root because it's a perfect square, \(\large\rm \sqrt4=?\)

OpenStudy (kellyspeakslouder):

2

zepdrix (zepdrix):

\[\large\rm f(g)=\sqrt{4(2x-1)}\]Ok good so it comes out as a 2,\[\large\rm f(g)=2\sqrt{(2x-1)}\]

OpenStudy (kellyspeakslouder):

yay!! thank you!! I have 4 other ones, but I don't think it's fair to keep you from helping everyone else.

zepdrix (zepdrix):

If you have other ones that you're stuck on, open them in a new thread. I'll try to stop by when I have some spare time :P Gotta eats mah sammich right now

zepdrix (zepdrix):

So it ended up being option A, right? Yayyy team

OpenStudy (kellyspeakslouder):

yay! and yes :)